OpenStudy (anonymous):

Square root of 5x +29 =x+3

OpenStudy (anonymous):



OpenStudy (anonymous):

square b s 5x+29=x^2+9+6x x^2+x-20=0 (x-4)(x+5)=0 x=4 or x=-5

OpenStudy (anonymous):

Thanks, I've always gotten confused when I worked problems that give me x=4 or x=-5.Can I put both numbers or one? My teacher didnt answer my question correctly.

OpenStudy (anonymous):

It's like x=this number or x= this...How do you determine which number to use?

OpenStudy (anonymous):

Both are valid answers,The solution should give 2 answers in a quadratic equation,Since the maximum power of x is 2

OpenStudy (anonymous):

I see...Thanks so much